# Mr-Ben: A Comprehensive Meta-Reasoning Benchmark for Large Language Models

### v1

@@ 6.1 Few Shot Prompting
As previously discussed and exemplified by our prompt template (Figure-10 in Appendix-C), our evaluation method is characterized by its high level of difficulty and complexity. In this experiment, we aimed to determine whether providing a few step-wise chain-of-thought (CoT) examples could improve model performance in terms of format adherence and reasoning quality. The results, as presented in Table-5 in Appendix, do not show a consistent pattern as the number of shots increases. While smaller language models like Gemma-2B exhibit performance improvements with additional shots, the performance of larger language models tends to fluctuate with an increasing number of shots. We hypothesize that for our complex tasks, the lengthy few-shot demonstrations may act more as a hindrance, providing distracting information rather than aiding in format adherence and reasoning. Our empirical findings suggest that a one-shot demonstration strikes the optimal balance between providing guidance and minimizing distraction. This supports our decision to focus on zero-shot versus one-shot comparisons in our primary experiments, as detailed in Table-2.
@@ 6.2 Self Refine Prompting
As suggested by [21], large language models typically cannot perform self-correction without external ground truth feedback. To explore whether this phenomenon occurs in our benchmark, we adopted a similar setting by prompting the language model to verify its own answer across a three-round interaction sequence: query, examine, and refine. Our prompting template, detailed in Figure-8 in Appendix-C, is minimalistic and designed solely to encourage the model to self-examine.
The results of this self-refinement process are recorded in Table-4. Notably, models smaller than Llama3-70B exhibit performance degradation with self-refinement, while larger models, such as GPT-4, show marginal benefits from the process. To understand the disproportionate improvement observed in the 70B model, we analyzed performance breakdown at the task level. These results are visualized in Figure-9 of Appendix-D.
This figure illustrates how the evaluated models modify their predictions on solution scoring during self-refinement. For instance, weaker models like Gemma-2B predominantly maintain incorrect predictions even after self-refinement. Conversely, from Llama3-8B to Llama3-70B, despite a significant portion of correct predictions shifting to incorrect ones, as previously reported by [21], our benchmark shows an increasing trend of incorrect predictions shifting to correct ones as model size increases. This shift results in the significant performance improvements observed in models like Llama3-70B.
@@ 6.3 Solution Correctness Prior
The results in Table-4 illustrate that the benefits of knowing the solution correctness prior generally increase with the model’s competence but begin to plateau at the level of sophisticated models like GPT-4. Specifically, the Gemma-2b model struggles significantly in our benchmark, showing nearly zero performance due to its limited ability to follow formats and comprehend complex tasks. Consequently, having the solution correctness prior does not improve its performance metrics. In contrast, models with moderate capabilities benefit substantially from this prior knowledge, which aids in accurately locating the first error step and elucidating the error reason. However, as model capabilities improve, the incremental benefits of this prior knowledge quickly diminish. For instance, GPT-4 shows only a marginal improvement in identifying the first error step and an almost negligible impact on error reason analysis when provided with the prior.
@@ 7 Conclusion
This paper highlights the importance of evaluating the reasoning capabilities of LLMs and proposes a comprehensive benchmark called Mr-Ben that addresses the limitations of existing evaluation methodologies. Mr-Ben consists of questions from a diverse range of subjects and incorporates a meta-reasoning paradigm, where LLMs act as teachers to evaluate the reasoning process. The evaluation of 15 LLMs on Mr-Ben reveals previously unknown limitations and weaknesses in their reasoning abilities. Improving LLMs’ reasoning abilities on complex and nuanced tasks remains an open research question, and we encourage future work to develop upon Mr-Ben. We provide a qualitative error analysis on the failure and success cases in Appendix E, in hopes of inspiring future investigations. The limitations and broader impact of the paper are discussed in Appendix A.1 and A.2.
@@ Appendix A Appendix
A.1 Limitations
A.2 Broader Impact
Positive Societal Impacts
Positive Societal Impacts
| Subject | Pearson Correlation | P-value |
| Medicine | 0.094 | 0.0072 |
| Physics | -0.061 | 0.111 |
| Biology | 0.009 | 0.783 |
| Chemistry | -0.127 | 0.00018 |
| Coding | -0.199 | 0.0021 |
| Logic | 0.0002 | 0.995 |
| Math | -0.115 | 0.00049 |
Appendix B More Discussion on Biases
To quantitatively assess the relationship between the length of solutions and their correctness, Pearson-Correlation-Coefficients were calculated and reported in Table-6 in the Appendix. The result suggests varying dynamics across disciplines regarding how solution length impacts the likelihood of correctness. For subjects such as coding, chemistry and math, longer solutions are less likely to be correct, which could suggest that complexity or elaboration in responses may lead to mistakes or incorrect reasoning. For medicine, despite being weak, there’s a tendency for longer solutions to be slightly more correct, possibly due to more detailed or thorough explanations being favorable. For the other subjects, length of solution does not appear to significantly affect correctness, indicating that other factors likely play a more dominant role in determining solution quality. The overall Pearson Coefficients analysis reflects the distinct nature of problem-solving in each field of our benchmark.
Appendix C Evaluation Prompt

### v2

@@ 6.1 Few Shot Prompting
As previously discussed and exemplified by our prompt template (Figure 10 in Appendix-D), our evaluation method is characterized by its high level of difficulty and complexity. In this experiment, we aimed to determine whether providing a few step-wise chain-of-thought (CoT) examples could improve model performance in terms of format adherence and reasoning quality. The results, as presented in Table 9 in Appendix, do not show a consistent pattern as the number of shots increases. While smaller language models like Gemma-2B exhibit performance improvements with additional shots, the performance of larger language models tends to fluctuate with an increasing number of shots. We hypothesize that for our complex tasks, the lengthy few-shot demonstrations may act more as a hindrance, providing distracting information rather than aiding in format adherence and reasoning. Our empirical findings suggest that a one-shot demonstration strikes the optimal balance between providing guidance and minimizing distraction. This supports our decision to focus on zero-shot versus one-shot comparisons in our primary experiments, as detailed in Table 2.
@@ 6.2 Self Refine Prompting
As suggested by [31], large language models typically cannot perform self-correction without external ground truth feedback. To explore whether this phenomenon occurs in our benchmark, we adopted a similar setting by prompting the language model to verify its own answer across a three-round interaction sequence: query, examine, and refine. Our prompting template, detailed in Figure 8 in Appendix D, is minimalistic and designed solely to encourage the model to self-examine.
The results of this self-refinement process are recorded in Table 4. Notably, models smaller than Llama3-70B exhibit performance degradation with self-refinement, while larger models, such as GPT-4, show marginal benefits from the process. Conversely, from Llama3-8B to Llama3-70B, despite a significant portion of correct predictions shifting to incorrect ones, as previously reported by [31], our benchmark shows an increasing trend of incorrect predictions shifting to correct ones as model size increases. This shift results in the significant performance improvements observed in models like Llama3-70B.
To understand the disproportionate improvement observed in the 70B model, we analyzed performance breakdown at the task level. These results are visualized and discussed in Figure 9 of Appendix E. In short, we believe the lack of consistency does not necessarily indicate a more robust or advanced reasoning ability, despite the increase of the evaluation results.
@@ 6.3 Solution Correctness Prior
The results in Table 4 illustrate that the benefits of knowing the solution correctness prior generally increase with the model’s competence but begin to plateau at the level of sophisticated models like GPT-4. Specifically, the Gemma-2b model struggles significantly in our benchmark, showing nearly zero performance due to its limited ability to follow formats and comprehend complex tasks. Consequently, having the solution correctness prior does not improve its performance metrics. In contrast, models with moderate capabilities benefit substantially from this prior knowledge, which aids in accurately locating the first error step and elucidating the error reason. However, as model capabilities improve, the incremental benefits of this prior knowledge quickly diminish. For instance, GPT-4 shows only a marginal improvement in identifying the first error step and an almost negligible impact on error reason analysis when provided with the prior.
@@ 7 Conclusion
This paper highlights the importance of evaluating the reasoning capabilities of LLMs with process-oriented design and presents a comprehensive benchmark called MR-Ben that addresses the limitations of existing evaluation methodologies. MR-Ben consists of questions from a diverse range of subjects and incorporates a meta-reasoning paradigm, where LLMs act as teachers to evaluate the reasoning process. Our evaluation of a diverse suite of LLMs on MR-Ben reveals several key limitations and weaknesses. Many models struggle with identifying and correcting errors within reasoning chains, demonstrating difficulty in performing system-2 style thinking—such as scrutinizing assumptions, calculations, and intermediate steps. Furthermore, even state-of-the-art models often fail to maintain consistency across reasoning paradigms, exposing gaps in their generalization abilities. Additionally, our findings emphasize the importance of searching and reflecting on the solution space during inference. Models like the o1 series showcase the potential of scaling test-time computation, where frequent retrospection and iterative search through multiple solution paths significantly enhance reasoning performance. Nevertheless, improving LLMs’ reasoning abilities on complex and nuanced tasks remains an open research question, and we encourage future work to develop upon MR-Ben.
8 Acknowledgement
This work was supported in part by the Research Grants Council under the Areas of Excellence scheme grant AoE/E-601/22-R.
@@ Appendix A Appendix
A.1 Limitations
A.2 Broader Impact
Positive Societal Impacts
Negative Societal Impacts
A.3 Additional Related Work
Improving Reasoning Abilities of LLMs
To enhance the reasoning capabilities of LLMs, prior research primarily focuses on specific prompting techniques [11]. Existing efforts include few-shot prompting with intermediate steps augmented demonstrations [66, 72, 69] or zero-shot prompting with specific instructions [36, 74]. Although these methods have shown promising results, their effectiveness is often constrained by their task-specific nature and the labour-intensive process of designing prompts, leading to inconsistent outcomes across different tasks [75, 80]. Another strategy to facilitate reasoning involves instruction tuning or knowledge distillation, which elicits reasoning paths from LLMs without explicit prompting [13, 49, 26, 44]. These approaches typically involve resource-intensive fine-tuning over LLMs and require a large set of examples annotated with CoT.
Learning From Feedback
Improving LLMs through learning from feedback has become a prevalent strategy, notably through reinforcement learning from human feedback, which seeks to align LLMs with human values by refining their outputs based on feedback [53, 8]. However, this method faces challenges such as high costs due to manual labor and a lack of real-time feedback capabilities [20]. An alternative strategy involves using self-correcting LLMs, which rely on automated feedback to iteratively adapt and understand the consequences of their actions without relying on humans. This feedback can be derived from outside sources such as other models [70, 45], tools [24, 29], knowledge bases [21, 76], evaluation metrics [34, 67] or generation logits [73].
Appendix B Robustness of MR-Score
| Model | Coding | Phy. | Bio. | Math | Med. | Chem. | Logic |
| gpt-4-turbo | 83/55 | 137/15 | 164/11 | 305/46 | 194/25 | 166/27 | 192/16 |
| deepseek_coder | 100/38 | 145/7 | 167/8 | 321/30 | 200/19 | 172/21 | 193/15 |
| Qwen2-72B | 99/39 | 142/10 | 167/8 | 312/39 | 195/24 | 172/21 | 200/8 |
| Coding | Phy. | Bio. | Med. | Chem. | Logic | Math | |
| Agreement Ratio | 7/8 | 12/13 | 21/21 | 12/12 | 15/17 | 15/16 | 10/13 |
Question: Does the ACC_reason metric’s dependency on the judgments of different LLMs or human evaluators lead to variability in scoring ?
Answer: We would like to argue that due to the careful design of our evaluation mechanism, the automatic scoring of error reasons is both robust and economically feasible:
Multiple annotators: During the annotation stage, we collected multiple annotations for the first error reasons and potential error rectification from different annotators who agreed on the solution correctness and the first error step.
Proxy Model Evaluation: Based on the ground truth annotations collected from various perspectives, the proxy language model (e.g., GPT-4-Turbo) then examines the error reasons provided by evaluating models. Given the question/solution pair and information regarding the first error step, error reasons, and rectification, the potential flaws of the error reasons provided by the evaluating models are easy to diagnose under contrast.
ACC_reason robustness: Table-5 shows the scores of error reasons sampled from our evaluation results. For the same set of error reasons collected in each subject, three different models made their predictions on correctness/incorrectness. We can clearly see the consistency of their predictions among the three models over questions in all subjects. Since the MR-Score is a weighted metric, the final score variability is less than 1 percent in total.
Human-Model Agreement Rate: As mentioned in 3, the agreement rate between manual annotations and the GPT-4 predictions over 100 samples randomly collected from all subjects is 92%. Below is the exact detail of our setup:
We randomly collected 100 data instances where the evaluating model correctly identified the solution correctness and the first error step across all subjects. We then manually examined whether the proxy scoring model (e.g., GPT-4-Turbo-2024-04-09) correctly scored the error reasons of the evaluating models. Table-6 is the detailed composition of the ratio in which the author agrees with the proxy scoring model. The annotation time varies significantly across subjects, as some problems—such as coding and chemistry—can take more than 10 minutes to evaluate, while subjects like biology are easier to assess. This high agreement rate further supports the reliability of our evaluation, thus avoiding the need for manual annotation of potentially 138,000 problems (6,000 benchmark size times 23 models evaluated).
| =0 | =1 | =0 | =1 | =0 | =1 | =0 | =1 | =0 | =1 | =0 | =1 | =0 | =1 | =0 | =1 | ||||||||
| MR-Scores | |||||||||||||||||||||||
| MCC-Matthews Correlation Coefficient | |||||||||||||||||||||||
| Claude3-Haiku | 13.96 | 17.72 | 16.47 | 13.62 | 15.09 | 10.74 | 16.54 | 19.96 | 8.52 | 8.35 | 6.21 | 4.94 | 4.36 | 0 | 11.59 | 10.76 | |||||||
| GPT-3.5-Turbo | 10.72 | 19.44 | 16.66 | 21.33 | 17.48 | 17.45 | 18.24 | 12.6 | 11.19 | 13.28 | 4.07 | 0 | 12.35 | 12.35 | 12.96 | 13.78 | |||||||
| Deepseek-Coder-33B | 7.51 | 8.57 | 11.73 | 6.81 | 9.69 | 21.06 | 9.98 | 7.94 | 1.62 | 6.28 | 0 | 0 | 26.18 | 15.44 | 9.53 | 9.44 | |||||||
| Deepseek-Coder-7B | 4.96 | 9.79 | 8.77 | 6.72 | 9.05 | 10.82 | 10.49 | 9.39 | 5.02 | 3.17 | 3.22 | 2.58 | 10.91 | 6.27 | 7.49 | 6.96 | |||||||
| LlaMA3-8B | 19.37 | 21.15 | 16.24 | 18.64 | 26.55 | 21.87 | 25.99 | 28.6 | 14.92 | 18.95 | 11.8 | 16.24 | 14.54 | 15.72 | 18.49 | 20.17 | |||||||
| Phi3-3.8B | 27.66 | 28.48 | 21.61 | 21.44 | 22.29 | 25.17 | 30.92 | 33.37 | 17.36 | 14.9 | 13.03 | 9.56 | 14.48 | 18.76 | 21.05 | 21.67 | |||||||
| Qwen1.5-72B | 33.64 | 42.44 | 31.4 | 31.56 | 29.2 | 23.28 | 35.47 | 36.47 | 21.76 | 29.64 | 24.42 | 27.74 | 13.8 | 15.69 | 27.1 | 29.55 | |||||||
| Deepseek-67B | 43.61 | 41.73 | 24.16 | 28.77 | 24.95 | 23.87 | 36.58 | 37.29 | 27.8 | 28.93 | 26.74 | 25.09 | 28.23 | 29.06 | 30.3 | 30.68 | |||||||
| LlaMA3-70B | 45.67 | 56.14 | 40.34 | 41.3 | 32.76 | 30.94 | 41.72 | 52.12 | 33.18 | 37.75 | 32.0 | 33.87 | 47.86 | 29.67 | 39.08 | 40.26 | |||||||
| Mistral-Large | 41.67 | 49.0 | 34.24 | 33.47 | 29.0 | 37.05 | 41.99 | 47.07 | 23.76 | 32.05 | 25.66 | 33.25 | 37.05 | 33.52 | 33.34 | 37.92 | |||||||
| Deepseek-v2-236B | 52.96 | 53.38 | 41.81 | 46.48 | 43.75 | 40.53 | 54.32 | 50.15 | 37.61 | 44.53 | 36.36 | 35.41 | 45.89 | 35.7 | 44.67 | 43.74 | |||||||
| GPT-4-Turbo | 63.33 | 62.59 | 52.9 | 52.7 | 50.67 | 52.84 | 53.05 | 54.59 | 56.79 | 54.66 | 40.95 | 42.94 | 52.5 | 57.53 | 52.88 | 53.98 | |||||||
| Accuracy of First Error Step | |||||||||||||||||||||||
| Claude3-Haiku | 2.15 | 3.1 | 1.4 | 1.12 | 2.38 | 1.59 | 1.77 | 4.42 | 1.69 | 0.68 | 1.01 | 0.29 | 0.0 | 0.0 | 1.49 | 1.6 | |||||||
| GPT-3.5-Turbo | 2.86 | 4.53 | 4.2 | 4.76 | 4.37 | 3.84 | 2.87 | 8.17 | 2.37 | 3.05 | 1.73 | 7.63 | 0.61 | 2.44 | 2.72 | 4.92 | |||||||
| Deepseek-Coder-33B | 14.83 | 10.29 | 14.94 | 12.36 | 14.69 | 10.92 | 15.67 | 16.31 | 14.54 | 12.43 | 12.22 | 18.18 | 5.49 | 3.05 | 13.2 | 11.93 | |||||||
| Deepseek-Coder-7B | 21.77 | 18.18 | 23.28 | 19.83 | 23.41 | 20.03 | 24.46 | 23.18 | 23.29 | 26.09 | 20.03 | 24.72 | 4.27 | 6.1 | 20.07 | 19.73 | |||||||
| LlaMA3-8B | 14.35 | 14.35 | 17.53 | 8.62 | 20.29 | 7.8 | 14.16 | 11.59 | 13.13 | 8.41 | 13.64 | 11.36 | 17.68 | 9.76 | 15.83 | 10.27 | |||||||
| Phi3-3.8B | 12.68 | 11.48 | 16.38 | 12.07 | 17.69 | 16.12 | 18.03 | 17.17 | 12.78 | 8.76 | 10.23 | 6.96 | 8.54 | 9.15 | 13.76 | 11.67 | |||||||
| Qwen1.5-72B | 11.48 | 15.31 | 10.63 | 11.49 | 10.79 | 9.88 | 12.45 | 14.38 | 11.03 | 13.49 | 8.1 | 10.94 | 1.83 | 4.27 | 9.47 | 11.39 | |||||||
| Deepseek-67B | 13.16 | 19.14 | 19.25 | 21.84 | 20.81 | 22.11 | 17.17 | 23.39 | 14.71 | 12.08 | 12.78 | 13.49 | 12.2 | 14.02 | 15.72 | 18.01 | |||||||
| LlaMA3-70B | 15.79 | 22.25 | 14.66 | 18.39 | 13.65 | 15.08 | 17.81 | 22.32 | 14.36 | 16.64 | 14.91 | 14.91 | 26.83 | 13.41 | 16.86 | 17.57 | |||||||
| Mistral-Large | 18.38 | 25.54 | 26.33 | 28.29 | 27.28 | 33.25 | 26.05 | 27.59 | 16.92 | 19.29 | 14.53 | 14.96 | 19.51 | 19.51 | 21.29 | 24.06 | |||||||
| Deepseek-v2-236B | 27.51 | 35.41 | 37.64 | 40.23 | 36.28 | 34.33 | 33.91 | 37.55 | 27.5 | 32.57 | 22.87 | 27.56 | 33.54 | 26.83 | 31.32 | 33.5 | |||||||
| GPT-4-Turbo | 41.77 | 46.06 | 42.58 | 46.5 | 46.49 | 47.81 | 42.6 | 46.8 | 37.06 | 36.72 | 29.93 | 33.24 | 50.61 | 59.15 | 41.58 | 45.18 | |||||||
| Accuracy of First Error Reason | |||||||||||||||||||||||
| Claude3-Haiku | 1.67 | 2.63 | 0.56 | 0.84 | 1.19 | 0.93 | 0.22 | 2.21 | 1.18 | 0.34 | 0.72 | 0.29 | 0.0 | 0.0 | 0.79 | 1.03 | |||||||
| GPT-3.5-Turbo | 1.19 | 2.63 | 2.24 | 1.96 | 1.85 | 1.46 | 0.88 | 3.53 | 1.35 | 1.69 | 0.72 | 4.17 | 0.61 | 1.83 | 1.26 | 2.47 | |||||||
| Deepseek-Coder-33B | 2.87 | 1.44 | 2.01 | 1.15 | 1.69 | 2.21 | 2.15 | 1.93 | 2.63 | 1.05 | 1.85 | 2.56 | 3.05 | 1.83 | 2.32 | 1.74 | |||||||
| Deepseek-Coder-7B | 5.98 | 5.02 | 6.03 | 4.6 | 5.98 | 7.93 | 5.79 | 6.22 | 4.9 | 5.08 | 6.25 | 5.54 | 3.05 | 5.49 | 5.43 | 5.7 | |||||||
| LlaMA3-8B | 7.66 | 6.7 | 4.89 | 2.3 | 7.28 | 4.55 | 6.22 | 7.08 | 4.73 | 3.33 | 5.97 | 5.97 | 15.24 | 7.93 | 7.43 | 5.41 | |||||||
| Phi3-3.8B | 8.13 | 6.7 | 6.9 | 5.75 | 7.02 | 6.37 | 9.66 | 10.52 | 5.78 | 5.08 | 5.4 | 2.7 | 7.32 | 7.32 | 7.17 | 6.35 | |||||||
| Qwen1.5-72B | 10.29 | 12.2 | 6.9 | 7.76 | 5.85 | 4.81 | 6.22 | 9.44 | 8.06 | 9.46 | 6.11 | 8.24 | 1.22 | 3.05 | 6.38 | 7.85 | |||||||
| Deepseek-67B | 8.85 | 11.24 | 8.62 | 10.06 | 8.32 | 9.49 | 7.73 | 12.23 | 9.46 | 5.6 | 8.81 | 10.37 | 10.37 | 10.37 | 8.88 | 9.91 | |||||||
| LlaMA3-70B | 13.16 | 18.42 | 9.77 | 13.51 | 8.58 | 10.27 | 11.59 | 15.88 | 10.68 | 13.49 | 10.94 | 11.08 | 24.39 | 13.41 | 12.73 | 13.72 | |||||||
| Mistral-Large | 15.27 | 21.0 | 19.05 | 20.45 | 15.1 | 21.72 | 16.56 | 18.54 | 13.03 | 14.21 | 11.22 | 11.94 | 17.07 | 17.68 | 15.33 | 17.94 | |||||||
| Deepseek-v2-236B | 22.25 | 31.58 | 25.57 | 30.17 | 25.1 | 23.28 | 22.96 | 28.11 | 21.54 | 27.5 | 18.89 | 24.15 | 29.88 | 23.78 | 23.74 | 26.94 | |||||||
| GPT-4-Turbo | 39.14 | 42.0 | 38.38 | 41.46 | 40.4 | 41.06 | 36.64 | 42.16 | 32.83 | 32.83 | 27.63 | 30.07 | 50.61 | 56.1 | 37.95 | 40.81 | |||||||
Question: Is the MR-Score sensitive to different weightings? Is MR-Score a robust unified metric?
Table-7 shows breakdown performance for models in all four metrics (MR-Score, MCC, ACC_step, and ACC_reason):
1.
Metric Robustness: Due to the progressive nature of the definitions of our subtasks (e.g., the success of subsequent tasks depends on the previous ones), we can see the diminishing trend in the scores of MCC, ACC_step, and ACC_reason. However, thanks to the design of our evaluation mechanism and metrics, the score rankings of different models stay in relatively stable order across metrics. In other words, we have not observed any model that excels in determining the solution correctness (thus high in MCC) but is unable to explain the rationale behind it (e.g., low in ACC_reason).
2.
Task Difficulties: As shown in the breakdown table, the ACC_reason metric is more discriminative than the MCC metric for competent models but vice versa for the less competent ones. This aligns with our intuition that generally more difficult questions are more discriminative for strong candidates, while weaker ones are simply incapable of solving them. This phenomenon could in part explain why in general the MR-Score is not very sensitive to minor changes in the weightings assigned to the subtasks, since the differentiability of the subtask metrics tends to reconcile with each other under different scenarios.
3.
Differentiability and Interpretability: The weights of the MR-Score are ultimately decided by considering both the discriminative ability and the interpretability. To best differentiate models with different evaluation results, we conducted a thorough grid search to investigate the impact of the weightings. Since the weightings calculated returned a few optimal instances, we deliberately selected the one that assigns higher scores to more difficult tasks. We believe the current weighting ratio strikes a good balance between interpretability and differentiation: For example, GPT-4-Turbo, Deepseek-v2-236B, and Mistral-Large achieve 86.4%, 78.5%, and 81.2% respectively in MMLU but score 43.2%, 29.4%, and 21.3% in our benchmark.
Appendix C More Discussion on Biases
To quantitatively assess the relationship between the length of solutions and their correctness, Pearson-Correlation-Coefficients were calculated and reported in Table-8 in the Appendix. The result suggests varying dynamics across disciplines regarding how solution length impacts the likelihood of correctness. For subjects such as coding, chemistry and math, longer solutions are less likely to be correct, which could suggest that complexity or elaboration in responses may lead to mistakes or incorrect reasoning. For medicine, despite being weak, there’s a tendency for longer solutions to be slightly more correct, possibly due to more detailed or thorough explanations being favorable. For the other subjects, length of solution does not appear to significantly affect correctness, indicating that other factors likely play a more dominant role in determining solution quality. The overall Pearson Coefficients analysis reflects the distinct nature of problem-solving in each field of our benchmark.
| Subject | Pearson Correlation | P-value |
| Medicine | 0.094 | 0.0072 |
| Physics | -0.061 | 0.111 |
| Biology | 0.009 | 0.783 |
| Chemistry | -0.127 | 0.00018 |
| Coding | -0.199 | 0.0021 |
| Logic | 0.0002 | 0.995 |
| Math | -0.115 | 0.00049 |
@@ Appendix C More Discussion on Biases
Appendix D Evaluation Prompt